# Learning to Infer Parameterized Representations of Plants from 3D Scans

### v1

@@ 3.2 Recursive Neural Networks
Recursive Neural Networks (RvNN) [12] are a class of deep neural networks that apply the same network architecture recursively on input that is structured as a tree graph to give a single representation in a latent space. The term recursive refers to the network architecture being applied to the output of step $i-1$ during step $i$. The recursivity of RvNNs allows to handle input of varying size, as the network is applied in a bottom-up fashion from the leaves to the root. These networks learn hierarchical dependencies and correlations between input components.
Socher et al. [46, 45] introduced an RvNN that learns information on different domains, including natural language processing and natural scene image processing, by recursively parsing a tree graph with binary structure. Models based on RvNN have been applied to natural language processing [9], 3D indoor scene generation [28, 17], blood vessel synthesis [16], and 3D shape structure recovery [35], segmentation [55] and generation [27].
Of particular interest for our work are recursive auto-encoders for binary trees, such as the ones used to generate 3D shapes of man-made objects [27]. A recursive auto-encoder learns on a binary tree structure, where all nodes can be represented in a latent space of dimension $dim_{\mathcal{S}}$. The encoder follows the tree structure and recursively merges pairs of inputs to form a new point in the same latent space until the full tree is represented as a single point in latent space. Inversely, the decoder recursively decodes a single point into two points in the same latent space until the full tree structure is decoded.
Figure 2 shows an example of the network architecture of a recursive encoder-decoder pair. The encoder-decoder pair allows to learn correlations between continuous properties encoded in features along the tree. An RvNN typically contains different types of encoder-decoder pairs. This allows to implement different operations, such as merging or splitting different parts.
During decoding, given a point in latent space, an algorithm is needed to decide which decoder to apply and when to stop the recursion. To achieve this, the encoders and decoders are trained with ground truth structural information. During encoding, the information of which encoder to apply is available. For decoding, the RvNN jointly learns a set of classifiers that is applied before each decoding step.
L-Strings have tree-like hierarchical structure, and we use RvNNs to encode them into latent points while capturing the relationships between the different modules, and their features’ correlations along the tree. We also use RvNNs in the opposite direction, to retrieve the tree structure and its parameters.
@@ 4 Method
4.1 Overview
We aim to obtain a tree-like parametric representation $l$ in the form of an L-String from an input point cloud $P$ that stems from an unstructured 3D scan of a plant. Learning a direct regression from the space of 3D point clouds to the space $\mathcal{L}$ of L-Strings is a difficult problem, as the point clouds have varying number of points and are unstructured, and as the L-Strings varies both in discrete (i.e. number and type of modules) and continuous (i.e. values of the parameters describing angles and lengths) ways.
To address this problem, we combine an RvNN learned on $\mathcal{L}$ with an encoder that maps a point cloud to $\mathcal{S}$, as shown in Fig. 3. During training, we first learn the latent space $\mathcal{S}$ that allows to represent L-Strings of a fixed plant species. This allows to represent 3D plants of varying shapes, with different number of organs, and different skeletal structures using latent representations $s\in\mathcal{S}$, and is explained in Sec. 4.2. Second, we train a neural network called point cloud encoder that maps an input point cloud $P$ to a point $s\in\mathcal{S}$, as outlined in Sec. 4.3). During inference, the input point cloud $P$ is encoded in the latent space as a point $s$ using the point cloud encoder, and subsequently decoded to an L-String using the recursive decoder, as detailed in Sec. 4.4.
@@ 4.2 Representing L-Strings in Latent Space
We aim to learn a latent space $\mathcal{S}$ that allows to represent various static instances of the same species of plant. Plants are represented in the space of L-Strings $\mathcal{L}$ and can be generated with a set of L-System rules. However, the branching structure of the L-Strings may not always correspond to a binary tree structure. We therefore first simplify the L-Strings by summarizing the information of modules that always occur together in the plant species. The resulting combined modules are called nodes in the following, and we design the combination rules to guarantee a binary tree structure after combination. To simplify notations, we call the simplified L-String $l\in\mathcal{L}$ in the following.
Our goal is to learn an encoder function $E:\mathcal{L}\to\mathcal{S}$, and a decoder function $D:\mathcal{S}\to\mathcal{L}$, such that $l\approx D(E(l)),\;\;\forall l\in\mathcal{L}$.
Learning to encode different shape and structural information in latent points of fixed dimension is challenging. Inspired by [27], we use an RvNN to learn the hierarchical relations between the modules of the L-String, leveraging the recursive nature of plant structures [20].
To achieve this, each node of the tree is represented individually in a point in the latent space, $s_{i}\in\mathcal{S}$, using a node auto-encoder. In a second step, each subtree of the L-String is represented in $\mathcal{S}$ using a recursive auto-encoder. This procedure is carried out recursively from the tree-graph leaves up to its root.
@@ 4.2.1 Node Auto-encoders
Each type of node has a different set of parameters, e.g. angles, widths, radii. This results in nodes that have different dimensionality in general, and that are not directly comparable. To allow nodes with different number of parameters to be used as input in an RvNN, we first map the information of each node to $\mathcal{S}$. To achieve this, a node encoder-decoder pair is learned for each node type. In the following, let $E_{node,i}$ and $D_{node,i}$ denote the encoder and decoder for each type of node, with $i=1,\ldots,N$ and $N$ the number of types of nodes. Both $E_{node,i}$ and $D_{node,i}$ consist of a fully connected linear layer, followed by a $tanh$ activation.
@@ 4.2.2 Recursive Auto-encoders
After applying node encoding, an input L-String is represented as a binary tree where all nodes are individually represented as points in $\mathcal{S}$. It can then serve as input to an RvNN auto-encoder. To recursively merge or split the nodes, we design encoder-decoder pairs based on the relationship of the input nodes in the tree. Two nodes to be merged can be either siblings or have a parent-child relationship in the tree.
Based on this observation, we consider the two merging/splitting criteria shown in Figure 4. For each criterion, we learn one recursive encoder-decoder pair. Both the sibling encoder-decoder $(E_{sib},D_{sib})$ and the parent-child encoder-decoder $(E_{pc},D_{pc})$ are implemented as shown in Fig 2.
@@ 4.2.3 Auxiliary Classifiers
In addition to the encoder-decoder pairs, it must be decided how to properly split the points to reconstruct the L-String structure. Each point in $\mathcal{S}$ is either a result of a merging process by one of the recursive encoders ($E_{sib}$, $E_{pc}$), or it represents an individual node. For the RvNN to decode structural information, it needs to learn to choose the appropriate decoder ($D_{sib}$, $D_{pc}$), or to end the recursive splitting. To predict the right decoder, we associate classes for the different splitting options (siblings, parent-child, stop) and jointly train a classifier on them.
Due to different node types, the network should needs to learn which node decoder $E_{node,i}$ to apply to reconstruct each node of the L-string. For that, a second classifier is trained, that allows to predict the node type for a latent point $s$ that is fully split. This classifier, called $C_{node}$ has $N$ output classes.
Both classifiers share the architecture of a two-layer fully connected neural network. The first layer maps the input latent point to a hidden representation using a linear transformation followed by a $tanh$ activation. The second layer maps the hidden features to an output vector that corresponds to the different classes.
@@ 4.2.4 Training
All auto-encoders are trained by minimizing a reconstruction loss at the node level. For node $n$ in L-String $l$, with a set of parameters, the reconstruction loss $L_{rec}(n)$ is the mean squared error between the input node parameter vector that contains the plant part information and its reconstruction. The reconstruction loss for the full L-String is
@@
At each recursive encoding step $i$, the output point $s$ is a result of merging two points ($s_{1}^{(i)}$, $s_{2}^{(i)}$). Then in the inverse decoding step, $s$ will be split into ($\hat{s}_{1}^{(i)},\hat{s}_{2}^{(i)}$) that are supposed to represent the same information as ($s_{1}^{(i)}$, $s_{2}^{(i)}$) and should fall in the same region in $\mathcal{S}$. To learn a structured representation space into which we can map information from another modality (point clouds), we encourage corresponding latent points during encoding and decoding to be located close-by in $\mathcal{S}$ using a step-based recursion loss $L_{step}$ that is defined as
$L_{step}=\sum_{i}\left(\left(s_{1}^{(i)}-\hat{s}_{1}^{(i)}\right)^{2}+\left(s_{2}^{(i)}-\hat{s}_{2}^{(i)}\right)^{2}\right).$ (2)
@@
$L_{total}=L_{rec}+L_{step}+L_{split}+L_{node}.$ (3)
@@ 4.3 Point Cloud Encoder
After representing L-Strings in latent space $\mathcal{S}$ has been learned, 3D plants can be represented as latent points $s\in\mathcal{S}$, and $s$ can be decoded into an L-String parametric representation $l$. To allow retrieving an L-String $l$ from an input point cloud $P$, we learn a new mapping function $E_{points}$ from the space of point clouds to $\mathcal{S}$ using a PointNet architecture [42].
To train $E_{points}$, we take advantage of paired input data, for which we are given both the point cloud $P$ and its corresponding L-String $l$. Passing $l$ through the recursive L-String encoder produces a latent point $s$. By applying $E_{points}$ on $P$, we obtain $\hat{s}$. The training optimizes the loss function
@@
where $j$ loops over all samples in the training set. This loss encourages the point cloud encoder function $E_{points}$ to represent $P$ at the location $\hat{s}\in\mathcal{S}$ that represents its corresponding L-String $s$.
@@ 4.4 Inferring L-Strings from Input Point Clouds
Errors in the predicted module parameters of the reconstructed L-String can have a propagating effect on the global geometry of the plant reconstruction. For example, errors in predicting the angle of a stem that is located in the bottom of the plant can lead to deviation in the plant growth axis along the main stem. To avoid such deviations, we align the reconstructed plant with the input in an optimization framework. We optimize on certain parameters of the stem modules starting from the ones at the bottom of the plant and going up. This optimization is done on the angle and length parameters for the Chenopodium album plants, minimizing the one-sided Chamfer Distance from the reconstructed stem to the input point cloud. We then optimize on the parameters of the individual leaf modules that define the leaf size and curvature. The latter optimization is applied on the leaves in no particular order, minimizing the single directional Chamfer Distance between the reconstructed plant and the input point cloud. This results in a parametric L-String representation $l$ that allows for various downstream phenotyping tasks. In this paper, we focus on the following tasks.
3D Reconstruction
can be solved directly by applying the geometric interpretation rules on $l$ to retrieve the virtual plant 3D scene.
Skeletonization
is solved by applying geometric interpretation rules drawing organ skeletons on $l$.
Segmentation
is solved by applying geometric interpretation rules on $l$ and by keeping the labels of the organ types. The labels are propagated to $P$ by assigning each point in $P$ the most frequently assigned label among its $k$ nearest neighbors in the annotated point cloud corresponding to $l$.
@@ 5 Dataset
To train and evaluate our method, we design and generate a synthetic dataset of L-String and point cloud pairs of the Chenopodium Album plant. First, we generate the L-Strings using LPy platform [2], with L-system production and geometric interpretation rules, that we optimize to generate realistic Chenopodium virtual plants. We define different time functions for the different plant parameters that guide the plant growth for a range of $[12,19]$ days to get Chenopodium plants in early stages. Then, we use the labeled points sampling method from [4], to obtain the corresponding point clouds.
@@
  • Stem: diameter, length, growing angle and phyllotaxis angle.
@@
The dataset contains plants of different shapes and structures, in a way that for each different structure there is a number of plants of different shapes. We ensure the balance between the different structures in the dataset by fixing the number of different plants for each structure, in order to avoid any bias towards certain structures. We generate plants of 33 different structures with 100 different plants for each structure, resulting in a dataset of 3300 pairs of L-Strings and point clouds.The dataset is split into training ($80\%$), validation($10\%$) and test($10\%$) sets. Figure 5 shows examples of generated plants from the dataset.
We evaluate our method on a test set of point clouds of synthetic plants, that have structures seen in the training set. Additionally, we evaluate on the same test set with simulated Gaussian noise, and on monocular depth images of the set, to show our models robustness to acquisition noise and incomplete inputs.
Finally we make first tests of our method on 5 scans of real Chenopodium Album plants from [33], to test its ability to generalize to real data.
Figure 5: Examples of Chenopodium Album plants from our dataset.
@@ 6 Evaluation
This section presents the evaluation of our method with respect to three common phenotyping applications, namely 3D reconstruction, 3D skeleton extraction, and segmentation. For each application, we outline an evaluation protocol and compare to strong baselines.
6.1 Implementation Details
We implemented our method using Pytorch on Quadro RTX 5000 GPU. We learn a latent space $\mathcal{S}$ of dimension $dim_{\mathcal{S}}=100$ with the recursive encoder-decoder pair architecture in Fig 2 of hidden layer size $h_{recur}=200$, and the classifiers with a hidden layer $h_{class}=200$. We optimize the weights of the network using Adam optimizer with a learning rate of 0.001, decaying by factor of 0.5 every 50 epochs. For the point cloud encoder we train a PointNet point regression network with Adam optimiser on a learning rate of 0.001. The weights of the layers of all the trained networks are initialized using Xavier uniform initialization.
Testset Method Accuracy Completeness Hausdorff Silhouette # Comp.
Synthetic SIREN $\mathbf{0.0035\pm 0.024}$ $\mathbf{0.001\pm 0.0004}$ $\mathbf{0.165\pm 0.204}$ $0.0395\pm 0.01$ $34.1\pm 19.2$
Ours $0.0148\pm 0.019$ $0.0186\pm 0.021$ $0.595\pm 0.215$ $\mathbf{0.0203\pm 0.008}$ $\mathbf{1}$
Noise SIREN $0.0156\pm 0.076$ $\mathbf{0.0019\pm 0.0006}$ $\mathbf{0.324\pm 0.327}$ $0.1425\pm 0.032$ $256.8\pm 69.4$
Ours $\mathbf{0.014\pm 0.017}$ $0.0182\pm 0.0186$ $0.598\pm 0.216$ $\mathbf{0.0204\pm 0.008}$ $\mathbf{1}$
Depth images SIREN $\mathbf{0.0098\pm 0.123}$ $\mathbf{0.0027\pm 0.0029}$ $\mathbf{0.428\pm 0.437}$ $0.0485\pm 0.03$ $41.2\pm 20.5$
Ours $0.0165\pm 0.02$ $0.021\pm 0.035$ $0.615\pm 0.22$ $\mathbf{0.0208\pm 0.008}$ $\mathbf{1}$
Table 2: Comparison to SIREN [44] for the 3D reconstruction of plants from point clouds on our test sets of synthetic scans, noisy point clouds and depth images. We compare on geometric distance measures (accuracy, completeness and Hausdorff), and on perceptual and qualitative measures (silhouette error and the number of connected components, denoted # Comp.).
Our method works on binary trees, while the L-String trees in our dataset mostly don’t have a binary tree structure. We summarize the L-Strings into binary trees, by combining the modules that always occur together as individual nodes with concatenated parameter sets. In particular, we combine the two cotyledon modules into one Cotyledons node with 4 parameters. Stems are always followed by a petiole and a leaf, therefore we combine them in one node called Stem that has 12 parameters. In case of a branching, the branch module is combined with the stem, petiole and leaf modules to form a Branch node of 14 parameters. Finally, the first stem module of the tree forms a node on its own called Root of 3 parameters. Note that we do not duplicate the symmetric parameters in the cotyledons (first two), which explains the number of resulting parameters. As for the Root node, we do not include the angle parameter since it is fixed across all the plants.
For the segmentation experiment where we use the $k$ nearest neighbors algorithm, we set $k=10$. As for the reconstruction experiments, when using SIREN we need to input the point normals along with the point cloud, for this we use a normal estimation PCA-based method, which fits local planes to each point’s neighborhood.
@@ 6.1 Implementation Details
We quantitatively evaluate our method for 3D reconstruction of plants from input point clouds, on our test sets using five evaluation metrics. The first three are the commonly used metrics for 3D reconstruction methods, namely accuracy, completeness, and Hausdorff distance. Accuracy and completeness are computed using the Chamfer distance between the ground truth point cloud and uniformly sampled points on the reconstructed surface. Accuracy measures how closely the reconstructed points lie to the ground truth, while completeness measures how well the reconstruction covers all parts of the ground truth’s geometry. Hausdorff distance measures the largest distances between the reconstructed and ground truth points, offering a worst-case assessment of the geometric alignment.
As these metrics are only based on 3D distances and do not account for the visual consistency of the reconstructions, we additionally consider the number of connected components that are reconstructed and silhouette error. The silhouette error is computed as the average pixel-wise absolute difference between the binary masks of the reconstructed and ground-truth meshes, rendered from multiple viewpoints. For each viewpoint, both meshes are rendered to produce binary images (silhouettes), where foreground pixels indicate visible parts of the shape. The error is then measured by taking the absolute value of the pixel-wise subtraction between the two silhouettes, averaged across all pixels and viewpoints (we use 400 viewpoints). This metric evaluates the visual similarity between the ground truth and the reconstructed mesh. Finally, the number of connected components measures global consistency and should be $1$ for our dataset of isolated plants.
Figure 6: Qualitative results for 3D reconstruction by our method and SIREN [44], on synthetic clean, noisy and partial point clouds.
@@
We compare our method to SIREN [44], an implicit neural representation method that uses sine activation functions to model continuous signals, including signed distance functions (SDFs). For 3D reconstruction, the surface can be extracted as the zero level set of the SDF using Marching Cubes. SIREN was identified as the best-performing method for reconstructing plant geometry from point clouds by Prasad et al. [38], due to its ability to capture fine geometric details.
Our method produces a complete smooth mesh that captures the plant visual shape and is robust with respect to input noise and missing input. Table 2 shows the results, where one can note that SIREN performs well in terms of the 3D distance-based error measures accuracy, completeness and Hausdorff distance. SIREN’s performance degrades significantly for depth images, and even more for noisy data. While our method has higher numerical error for these measurements than SIREN, it is much more robust to partial observations and noise. When considering silhouette loss, our method significantly outperforms SIREN as the visual similarity of the results is much higher for our method, (see Fig. 6). While the number of connected components of the reconstruction is always $1$ for our method by construction, SIREN reconstructs highly fragmented shapes.
@@
Fig. 6 shows examples of SIREN’s and our reconstructions for the three types of test examples. Note that while our reconstructions may have some misalignments, they are well structured and similar to the ground truth. In contrast, SIREN reconstructs fragmented shapes close to the input point clouds that are globally dissimilar from the ground truth plants.

### v2

@@ 3.2 Recursive Neural Networks
Recursive Neural Networks (RvNN) [costa2003towards] are deep neural networks that apply the same network architecture recursively on structured input. The term recursive refers to the network being applied to the output of step i1i-1 during step $i$. The recursivity of RvNNs allows to handle input of varying size, as the network is applied bottom-up from the leaves to the root.
RvNNs have been applied to different domains [socher2010learning, socher2011parsing] including natural language processing [chuan-an-etal-2018-unified], 3D generation [li2019grains, 10018465, 10.1145/3072959.3073637], blood vessel synthesis [10.1007/978-3-031-43907-0_7], 3D shape structure recovery [Niu_2018_CVPR], and segmentation [Yu_2019_CVPR].
Of particular interest for our work are recursive auto-encoders for binary trees [10.1145/3072959.3073637]. A recursive auto-encoder learns on a binary tree structure, where all nodes can be represented in a latent space of dimension $dim_{\mathcal{S}}$. The encoder follows the tree structure and recursively merges pairs of inputs to form a new point in the same latent space until the full tree is represented as a single latent point. Inversely, the decoder recursively decodes a single point into two points in the same latent space until the full tree structure is decoded.
@@ 4 Method
Our goal is to infer an L-String $l$ from an input 3D scan of a plant represented as point cloud $P$. Learning a direct regression from the space of 3D point clouds to the space $\mathcal{L}$ of L-Strings is difficult, as point clouds have varying numbers of points and are unstructured, and as L-Strings vary both in discrete (i.e. number and type of modules) and continuous (i.e. values of the angle and length parameters) ways.
To address this problem, we combine an RvNN learned on $\mathcal{L}$ with an encoder that maps a point cloud to a latent space $\mathcal{S}$, as shown in Fig. 2. During training, we first learn the latent space $\mathcal{S}$ that allows to represent L-Strings of a fixed plant species, see Sec. 4.1. Second, we train a neural network called point cloud encoder that maps an input point cloud $P$ to a point $s\in\mathcal{S}$, see Sec. 4.2. During inference, the input point cloud $P$ is encoded in $s\in\mathcal{S}$ using the point cloud encoder, and subsequently decoded to an L-String using the recursive decoder, see Sec. 4.3.
@@ 4.1 Representing L-Strings in Latent Space
We aim to learn a latent space $\mathcal{S}$ that represents instances of a plant species. The branching structure of the L-Strings may not always correspond to a binary tree structure. We therefore first simplify the L-Strings by summarizing the information of modules that always occur together in the plant species. The definition of co-occurring modules is manually done once per species. The resulting combined modules are called nodes in the following, and we design the combination rules to guarantee a binary tree structure after combination. To simplify notations, we call the L-String with binary tree graph structure $l\in\mathcal{L}$ in the following.
Our goal is to learn an encoder function $E:\mathcal{L}\to\mathcal{S}$, and a decoder function $D:\mathcal{S}\to\mathcal{L}$, such that $l\approx D(E(l)),\;\;\forall l\in\mathcal{L}$. Learning to encode different shape and structural information in latent points of fixed dimension is challenging. Inspired by [10.1145/3072959.3073637], we use an RvNN to learn the hierarchical relations between the modules of the L-String, leveraging the recursive nature of plant structures [Godin.20103zn].
To achieve this, each node of the tree graph is represented individually as point in latent space, $s_{i}\in\mathcal{S}$, using a node auto-encoder. In a second step, each subtree of $l$ is represented in $\mathcal{S}$ using a recursive auto-encoder. This procedure is carried out recursively from the leaves to the root resulting in a latent point that represents the full $l$ in $\mathcal{S}$.
@@ 4.1.1 Node Auto-encoders
Each type of node has a different set of parameters, e.g. angles, widths, radii. This results in nodes that have different dimensionality in general, and that are not directly comparable. To allow nodes with different numbers of parameters to be used as input in an RvNN, we first map the information of each node to $\mathcal{S}$. To achieve this, a node encoder-decoder pair is learned for each node type. In the following, let $E_{node,i}$ and $D_{node,i}$ denote the encoder and decoder for each type of node, with $i=1,\ldots,N$ and $N$ the number of types of nodes. Both $E_{node,i}$ and $D_{node,i}$ consist of one fully connected linear layer, followed by a tanhtanh activation.
@@ 4.1.2 Recursive Auto-encoders
After applying node encoding, an L-String is represented as a binary tree where all nodes are individually represented as points in $\mathcal{S}$. It can then serve as input to an RvNN auto-encoder. To recursively merge or split nodes, we design encoder-decoder pairs based on the relationship of the input nodes in the tree. Two nodes to be merged can either be siblings or have a parent-child relationship in the tree.
Based on this observation, we consider the two merging/splitting criteria shown in Fig. 3. For each criterion, we learn one recursive encoder-decoder pair. Both the sibling encoder-decoder $(E_{sib},D_{sib})$ and the parent-child encoder-decoder $(E_{pc},D_{pc})$ are implemented as shown in Fig 4.
@@ 4.1.3 Auxiliary Classifiers
In addition to the encoder-decoder pairs, it must be decided how to properly split the points to reconstruct the L-String structure. Each point in $\mathcal{S}$ is either a result of a merging process by one of the recursive encoders ($E_{sib}$, $E_{pc}$), or it represents an individual node. For the RvNN to decode structural information, it needs to learn to choose the appropriate decoder ($D_{sib}$, $D_{pc}$), or to end the recursive splitting. To predict the right decoder, we associate classes for the different splitting options (siblings, parent-child, stop) and jointly train a classifier $C_{split}$ on them.
The network needs to learn which node decoder $E_{node,i}$ to apply to reconstruct each node of the L-String. For that, a second classifier is trained, that allows to predict the node type for a latent point $s$ that is fully split. This classifier is called $C_{node}$. Both classifiers use a two fully connected layers, with tanhtanh activation after the first layer.
@@ 4.1.4 Training
All auto-encoders are trained by minimizing a reconstruction loss at the node level. For node $n$ in L-String $l$, with a set of parameters, the reconstruction loss $L_{rec}(n)$ is the mean squared error between the input node parameter vector that contains the plant part information and its reconstruction. The parameters’ weights are different in the loss. The reconstruction loss for the full L-String is
@@
$L_{total}=L_{rec}+L_{split}+L_{node}.$ (2)
@@ 4.2 Point Cloud Encoder
After learning latent shape space $\mathcal{S}$, 3D plants can be represented as $s\in\mathcal{S}$, and $s$ can be decoded into a parametric L-String representation $l$. To infer an L-String $l$ from an input point cloud $P$, we learn a mapping function $E_{points}$ from the space of point clouds to $\mathcal{S}$ using a PointNet [DBLP:journals/corr/QiSMG16].
To train $E_{points}$, we take advantage of paired input data, containing both the point cloud $P$ and its corresponding L-String $l$. Passing $l$ through the recursive L-String encoder produces a latent point $s$. By applying $E_{points}$ on $P$, we obtain $\hat{s}$. The training optimizes the loss
@@
where $j$ loops over all training samples. This loss encourages the point cloud encoder $E_{points}$ to represent $P$ at the location $\hat{s}\in\mathcal{S}$ that represents its corresponding L-String.
@@ 4.3 Inferring L-Strings from Input Point Clouds
Errors in the predicted module parameters of the reconstructed L-String can lead to cumulated errors on the plant reconstruction. For example, errors in predicting the angle of a stem that is located in the bottom of the plant can lead to deviation in the plant growth axis along the main stem. To avoid such deviations, we align the reconstructed plant with the input in a test-time optimization framework. We optimize on parameters of the main stem modules starting from the ones at the bottom of the plant and going up. This optimization is done on 3D angle and length parameters. We then optimize on the parameters of the petiole modules that define the length and elasticity, and then on the parameters of leaf modules that define the leaf size and curvature. All the modules are optimized w.r.t. the bidirectional Chamfer Distance between the reconstructed plant and the input point cloud, and two optimization iterations from bottom to top are performed. This results in a parametric L-String representation $l$ that allows for various downstream phenotyping tasks. In this paper, we focus on the following tasks.
3D Reconstruction can be solved by applying the geometric interpretation rules on $l$ to retrieve the 3D plant.
Skeletonization is solved by applying geometric interpretation rules on $l$ to reconstruct all stems and optionally main veins of leaves with minimal width.
Segmentation is solved by applying geometric interpretation rules on $l$ and keeping the labels of the organ types. The labels are propagated to $P$ by assigning each point in $P$ the most frequently assigned label among its $k$ nearest neighbors in the annotated point cloud corresponding to $l$.
@@ 5 Dataset
To train and evaluate our method, we design and generate a synthetic dataset of corresponding L-String and point cloud pairs of the Chenopodium Album plant. First, we generate the L-Strings using L-Py platform [boudon2012py], with L-system production and geometric interpretation rules, that we optimize to generate realistic Chenopodium virtual plants. We defined different time functions for the different plant parameters that guide the plant growth for a range of $[8,14]$ days to get Chenopodium plants in early stages. Then, we used the labeled points sampling method from [chaudhury20203d] to obtain the corresponding point clouds.
@@
  • Stem: diameter, length, growing angle, bending angle and phyllotaxis angle.
@@
The dataset contains plants of different shapes and structures. We balance the different structures in the dataset by fixing the number of different plants for each structure. We generated plants of 10 different structures with 100 different plants per structure, resulting in a dataset of 1000 pairs of L-Strings and point clouds. The dataset is split into training ($80\%$), validation ($10\%$) and test ($10\%$) sets.
This dataset of realistic Chenopodium Album plants with ground truth parametric representation can serve as benchmark for plant reconstruction, segmentation and skeletonization tasks.
We evaluate on the test set with clean point clouds, point clouds with simulated Gaussian noise, and monocular depth images of the set, to show our model’s robustness. To test generalization to real data, we test our method on 5 scans of real Chenopodium Album plants from Mirande et al[mirande_2022_6962994].
@@ 6 Evaluation
In this section we evaluate our method for the three common phenotyping applications 3D reconstruction, 3D skeleton extraction, and segmentation. For each application, we outline an evaluation protocol and compare to strong baselines. All methods are run on a Quadro RTX 5000 GPU. Implementation details are in appendix.
Accuracy Completeness Size Time # Comp. LN Accuracy LAI Accuracy Topology
Clean
SIREN $\mathbf{0.0012}$ $\mathbf{0.0006}$ $780.88$ KB 7min 14s7\,min\,14\,s $33$
Ours $0.0059$ $0.0090$ $\mathbf{17.56}$ KB $\mathbf{4\,min\,1\,s}$ $\mathbf{1}$ $\mathbf{98\%}$ $\mathbf{93\%}$ $\mathbf{75\%}$
Noisy
SIREN $0.0121$ $\mathbf{0.0008}$ $780.88$ KB 7min 27s7\,min\,27\,s $268$
Ours $\mathbf{0.0054}$ $0.0074$ $\mathbf{17.56}$ KB $\mathbf{3\,min\,56\,s}$ $\mathbf{1}$ $\mathbf{98\%}$ $\mathbf{93\%}$ $\mathbf{78\%}$
Depth maps
SIREN $\mathbf{0.0013}$ $\mathbf{0.0022}$ $780.88$ KB 6min 44s6\,min\,44\,s $49$
Ours $0.0060$ $0.0089$ $\mathbf{17.57}$ KB $\mathbf{3\,min\,47\,s}$ $\mathbf{1}$ $\mathbf{98\%}$ $\mathbf{94\%}$ $\mathbf{78\%}$
Table 2: Comparison to SIREN [sitzmann2020implicitneuralrepresentationsperiodic] for 3D reconstruction on clean and noisy point clouds and depth maps. We compare geometric measures (accuracy, completeness), the model’s size and inference time, number of connected components of output (# Comp.), and phenotyping measures leaf number (LN) accuracy, leaf area index (LAI) accuracy, and topology. ✗: the method cannot output the information.
@@
We quantitatively evaluate our method using eight complementary evaluation measures. The first two are accuracy and completeness, which are commonly used metrics assessing geometric alignment. Accuracy is the unidirectional Chamfer distance from the reconstructed surface to the ground truth, and completeness is the unidirectional Chamfer distance from the ground truth to the reconstructed surface. The next two measures are the output representation’s size and the inference’s running time, which measure the efficiency and compactness of the methods. We further report the number of connected components of the output meshes. We also assess the ability of models to predict correctly the number of leaves, defined as the mean percentage of matching leaf count between the reconstruction and the ground truth, and the mean percentage accuracy of leaf area index (LAI), defined as the one-sided leaf area per unit ground surface area [watson_lai]. These two measures have been selected as they are commonly used in agronomy to calibrate crop models. Finally, topology accuracy is the percentage of reconstructed plants whose topology is correct (measured using a tree-edit distance [ferraro-hal-00827474]).
Figure 5: Comparison to SIREN [sitzmann2020implicitneuralrepresentationsperiodic] for 3D reconstruction.
@@
We compare our method to SIREN [sitzmann2020implicitneuralrepresentationsperiodic], an implicit neural representation method that uses sine activation functions to model continuous signals, including signed distance functions (SDFs). For 3D reconstruction, the surface is extracted as the zero level set of the SDF using Marching Cubes. SIREN was identified as the best-performing method to reconstruct plant geometry from point clouds by Prasad et al[prasad2022deep].
Table 2 shows the results. While SIREN performs well in terms of the 3D distance-based error measures accuracy and completeness for clean data, SIREN’s performance degrades significantly for noisy data. Our method is more robust w.r.t. noise and missing data, and outperforms SIREN in case of noise. Our representation is one to two orders of magnitude more compact than SIREN’s, and inference is twice as efficient. Unlike SIREN, which reconstructs a mesh without annotations, our method further allows to measure leaf number accuracy, LAI error and topology, and achieves very high accuracy on all these measures.
@@
Fig. 5 shows 3D reconstructions for the three types of test examples. Note that our reconstructions are smooth, similar to the ground truth, and robust w.r.t. input noise and missing data. In contrast, due to its computational strategy, SIREN reconstructs fragmented shapes close to the input point clouds that are globally dissimilar from the ground truth, especially for noisy input. Finally, Fig. 6 (second column) shows the 3D reconstructions obtained when applying our method to scans of real plants. Interestingly, despite local discrepancies, all plants are well reconstructed overall, which demonstrates that the methods generalizes well from virtual to real plant architecture data.
Figure 6: Our method’s results on scans of real plants [mirande_2022_6962994]. Scans are reconstructed and segmented well overall, and the extracted skeletons closely follow the plants’ topologies.